OpenStudy (anonymous):

HOW DO YOU WRITE 4% decimal

OpenStudy (skullpatrol):

you need to understand the percent equation

OpenStudy (anonymous):

4% means 4/100 = 0.04

OpenStudy (igreen):

4 / 100 = ? @noterdame12345

OpenStudy (igreen):

@Karolyn101 Don't give answers.

OpenStudy (skullpatrol):

It defines what a percent is

OpenStudy (anonymous):

0.04

OpenStudy (igreen):

Yep, so 0.04 is your answer. @noterdame12345

OpenStudy (anonymous):

oopppppppppppppppppppppps ps

OpenStudy (skullpatrol):

yes, do you want the defining equation?

OpenStudy (anonymous):

(\_/) (-.-) (> <)

OpenStudy (chosenmatt):

yep so it's 0.04

OpenStudy (morganchavez):

0.04

OpenStudy (morganchavez):

yup

OpenStudy (skullpatrol):

Percent means "hundredths" or "divided by 100" @noterdame12345 do you understand what that means in equation form? $$\Huge \frac{x}{100}=\frac{a}{b}$$

OpenStudy (skullpatrol):

To understand all percent problems you must understand what the a, b, and x in this equation mean.

OpenStudy (skullpatrol):

Would you like to learn @noterdame12345 ?